# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v1

@@ 4.1. Datasets
Cityscapes. The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types, i.e. person, rider, car, truck, bus, train, motorcycle, and bicycle. We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
@@
For Cityscapes-C, both short-term and long-term CTTA tasks are explored. Our short-term CTTA task selectively focuses on the latter 12 corruptions, i.e. Defocus Blur, Frosted Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic, Pixelate, and JPEG. For the long-term CTTA task, we prioritize the five corruptions with direct relevance to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus. To mimic the scenario in real life where similar environments might be revisited, and to evaluate the forgetting effect of our methods in the long term, we repeat the adaptation to the target domain group 10 times.
@@
ACDC. The ACDC dataset (Sakaridis et al., 2021), namely the adverse conditions dataset, shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022; Gan et al., 2023), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images from each condition. Similarly, we adapt the source-trained model to continually to the target domain groups 10 cycles for the long-term CTTA task.
@@ 4.2. Implementation Details
We adopt the Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the object detection model is required to adapt toward a continuous influx of images. The source and student models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More implementation details are provided in the supplementary material.
@@ 4.3. Baselines and Compared Approaches
To establish the efficacy of our proposed CTAOD, we compare it with five baselines across various setting types for a fair comparison, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), Tent (Wang et al., 2020), and CoTTA (Wang et al., 2022). Specifically, ”Source” represents the source model, i.e., Faster R-CNN (Ren et al., 2015), solely pre-trained on the source domain. Moreover, Memclr (VS et al., 2023b) and IRG (Ren et al., 2015) are SFDA object detection methods. MemCLR (VS et al., 2023b) integrates cross-attention with contrastive learning (CL), while IRG (VS et al., 2023a) incorporates instance relation graph and CL. Furthermore, Tent (Wang et al., 2020) updates the affine parameters through entropy minimization in TTA. CoTTA (Wang et al., 2022) utilizes weight-averaged predictions and random recovery neurons to tackle CTTA, while SVDP (Yang et al., 2024) explores sparse visual prompts for CTTA dense prediction.
Table 4. Experimental results (mAP@0.5) of SHIFT short-term CTTA task. We evaluate the adaptation performance by continually adapting the source model to the four conditions.
Time $t\xrightarrow{\hskip 195.12767pt}$ All
Condition Cloudy Overcast Rainy Foggy Mean Gain
Source 51.8 41.5 43.8 33.9 42.7 /
Tent 50.9 39.5 36.3 23.1 37.5 -5.2
CoTTA 51.1 40.1 40.7 29.9 40.5 -2.2
SVDP 52.0 41.0 43.8 35.9 43.2 +0.5
IRG 51.9 40.6 42.7 34.3 42.4 -0.3
MemCLR 51.8 40.4 42.6 35 42.4 -0.3
Ours 52.3 41.3 44.1 36.8 43.6 +0.9
@@ 4.4. Results and Analysis
Table 5. Experimental results (mAP@0.5) of Cityscapes-to-ACDC long-term CTTA task. We evaluate the long-term adaptation performance by continually adapting the source model to the four conditions ten times. To save space, we display selected rounds of results, and full results are provided in the supplementary material.
@@
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean Gain
Source (Ren et al., 2015) 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36.0 /
Tent (Wang et al., 2020) 52.4 18.6 33.4 38.9 51.7 17.4 31.4 36.0 45.8 14.6 28.1 28.5 35.0 9.5 21.5 18.3 30.5 -5.5
CoTTA (Wang et al., 2022) 53.7 19.7 38.0 42.4 53.1 19.7 37.7 42.9 51.3 19.0 36.5 41.8 50.9 19.1 36.0 42.4 37.8 +1.8
SVDP (Yang et al., 2024) 52.8 20.0 35.6 42.0 54.6 23.5 38.7 43.8 52.8 24.0 38.6 43.9 51.8 23.6 38.2 43.0 39.5 +3.5
IRG (VS et al., 2023a) 52.7 20.6 36.0 42.9 53.6 23.2 38.1 45.6 51.0 23.1 37.1 42.8 49.7 22.5 36.5 40.7 38.9 +2.9
MemCLR (VS et al., 2023b) 52.9 21.2 35.2 42.8 52.8 23.2 38.2 44.4 51.4 23.2 37.1 42.9 49.9 22.9 36.6 41.0 38.7 +2.7
Ours 52.9 20.4 34.5 42.9 54.5 23.8 39.6 45.8 53.5 25.1 39.0 45.3 52.5 24.3 38.8 44.7 40.3 +4.3
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of the proposed method on the synthetic continual shift dataset (i.e. Cityscapes-to-Cityscapes-C adaptation tasks) in the short term. As depicted in Table 2, Tent (Wang et al., 2020) undergoes a slight decline in performance, dropping from 15.1 to 14.2 mAP relative to the source model. This downturn may be attributed to its dependency on a large batch size to update the parameters of the batchnorm layer, making it suboptimal for online adaptation. In contrast, CoTTA (Wang et al., 2022), IRG (VS et al., 2023a), and Memclr (VS et al., 2023b) exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which employs weight-averaged predictions. However, SVDP that utilizes visual prompts achieves the sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.6 mAP, consistently outperforming the above approaches. CTAOD ensures a more reliable adaptation to target domains characterized by intense and continual changes in the short term with high-quality pseudo-labels and feature representation.
As presented in Table 9, the long-term task outcomes reveal the source model’s relatively poor performance, with an average mAP of 16.4. Tent’s performance is markedly deteriorative across the timeline. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP that employ sparse visual prompts raise the performance to 25.3 mAP. Particularly, CTAOD yields a remarkable 12.6 mAP enhancement over the Source and surpasses all comparative baselines, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable adaptation amidst synthetic continual distribution shifts, across both short-term and long-term adaptation scenarios.
@@ 4.4.1. Synthetic Continual Distribution Shift
Figure 3. Qualitative results. We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to brightness (row 1) and defocus (row 2) corruption on the long-term Cityscapes-to-Cityscapes-C task.
@@ 4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset (i.e. SHIFT and Cityscapes-to-ACDC adaptation tasks). For the SHIFT short-term CTTA task, where each domain includes more complex images ranging from daytime to night, the results is shown in Table 4. While other baseline method suffer from performance decline, SVDP that employs fixed threshold for pseudo-labeling and prompts learning achieve a 0.5 mAP improvements. Moreover, CTAOD that introduces a DT strategy shows a superior performance of 43.6 mAP.
The experimental results of ACDC long-term CTTA task are summarized in Table 5. Tent, CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, CTAOD achieves a performance of 40.3 mAP, yielding an absolute improvement of 0.8 mAP over baselines. While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. This outcome underscores the capability of our method to foster robust adaptation toward the real-world continual distribution shift in the long term.
@@ 4.4.3. Qualitative Results
As shown in Figure 3, we provide some examples of detection results. We selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness and defocus corruption on the long-term Cityscapes-to-Cityscapes-C tasks. Compared with Source and MemCLR, CTAOD guides the model to learn better feature representations while effectively mitigating forgetting in the long-term adaptation. Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them. More visualization analysis is presented in the supplementary material.
@@
Mean-Teacher OCL DT DSR Mean Gain
1 $\checkmark$ 25.9 /
2 $\checkmark$ $\checkmark$ 27.4 +1.5
3 $\checkmark$ $\checkmark$ 27.1 +1.2
4 $\checkmark$ $\checkmark$ 26.7 +0.8
5 $\checkmark$ $\checkmark$ $\checkmark$ 28.5 +12.1
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.0 +12.6
@@ 4.4.4. Ablation Study
We conduct an ablation study to empirically assess the impact of the proposed OCL, DT, and DSR. As shown in Table 6, the integration of the mean-teacher architecture, which incorporates weak-strong augmentation and KL divergence distillation, elevates the performance to 25.9 mAP. Building upon the mean-teacher foundation, the proposed OCL, DT, and DSR achieve an additional 1.5, 1.2, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and DT modules raises the performance to 28.5 mAP. Furthermore, the inclusion of the DSR mechanism brings a 0.5 mAP boost in performance, culminating in a comprehensive model performance of 29.0 mAP
@@ Appendix A Appendix Overview
In this supplementary material, we provide additional details and analysis. In Section B, we provide implementation details on CTAOD and other baseline methodologies. In Section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research. In Section D, we discuss the limitations and prospective directions of our research.
@@ Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
@@ B.1. Implememtation of CTAOD
Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 8. Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1. The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Dynamic Threshold, and Data-driven Stochastic Restoration.
@@ B.2. Implememtation of Baseline Methods
The Tent (Wang et al., 2020), CoTTA (Wang et al., 2022), and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, the Tent approach is implemented by updating the parameters of the batch normalization in Faster R-CNN through entropy minimization. Meanwhile, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.

### v3

@@ 4.1. Datasets
Cityscapes. The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types. We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
@@
Our short-term CTTA task selectively focuses on the latter 12 corruptions. For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain groups following (Gan et al., 2023), namely Fog, Motion, Snow, Brightness, and Defocus. We repeat the adaptation to the target domain group 10 times to evaluate long-term performance.
@@
ACDC. The ACDC dataset (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition. Similarly, the source-trained model is continually adapted to the target domain group for 10 cycles.
@@ 4.2. Implementation Details
We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1, thereby emulating a real-world application scenario where the detection model adapts toward a continuous influx of images. The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More hyper-parameters and implementation details are provided in the appendix.
@@ 4.3. Baselines and Compared Approaches
We compare CTAOD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and CTAOD-upstop. Specifically, ”Source” represents the source model without adaptation. Memclr and IRG are SFDA object detection methods, also referred to as TTA methods (Wang et al., 2022). MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL. CoTTA utilizes weight-averaged predictions and random neurons recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction. “CTAOD-unstop” represents CTAOD without dynamic skipping.
@@ 4.4. Results and Analysis
Condition Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Fog Night Rain Snow Mean Gain Iter.
Source 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 52.3 18.7 33.5 39.6 36.0 / /
CoTTA 53.7 19.7 38.0 42.4 53.1 19.7 37.7 42.9 51.3 19.0 36.5 41.8 50.9 19.1 36.0 42.4 37.8 +1.8 16k
SVDP 52.8 20.0 35.6 42.0 54.6 23.5 38.7 43.8 52.8 24.0 38.6 43.9 51.8 23.6 38.2 43.0 39.5 +3.5 16k
IRG 52.7 20.6 36.0 42.9 53.6 23.2 38.1 45.6 51.0 23.1 37.1 42.8 49.7 22.5 36.5 40.7 38.9 +2.9 16k
MemCLR 52.9 21.2 35.2 42.8 52.8 23.2 38.2 44.4 51.4 23.2 37.1 42.9 49.9 22.9 36.6 41.0 38.7 +2.7 16k
Ours-unstop 52.9 20.4 34.5 42.9 54.5 23.8 39.6 45.8 53.5 25.1 39.0 45.3 52.5 24.3 38.8 44.7 40.3 +4.3 16k
Ours 52.8 20.1 35.2 42.9 54.4 24.6 38.6 45.5 53.5 25.6 38.5 46.2 53.5 25.4 37.4 45.3 40.4 +4.4 13.2k
Table 3. Experimental results (mAP@0.5) and adaptation iterations of Cityscapes-to-ACDC long-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions ten times. To save space, we display selected rounds of results, and full results are provided in the appendix. “-unstop” denotes no skipping.
@@
Figure 3. Qualitative results. We compare the detection results of the Source model, MemCLR, and CTAOD in the $10^{th}$ round of adaption to defocus corruption on the long-term Cityscapes-to-Cityscapes-C task.
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of CTAOD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks. As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate $\alpha$ for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions. Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches. Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to CTAOD-unstop. CTAOD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP. Particularly, CTAOD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the DSR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
Time $t\xrightarrow{\hskip 195.12767pt}$ All
Condition Cloudy Overcast Rainy Foggy Mean Gain Iter.
Source 51.8 41.5 43.8 33.9 42.7 / /
CoTTA 51.1 40.1 40.7 29.9 40.5 -2.2 19.6k
SVDP 52.0 41.0 43.8 35.9 43.2 +0.5 19.6k
IRG 51.9 40.6 42.7 34.3 42.4 -0.3 19.6k
MemCLR 51.8 40.4 42.6 35 42.4 -0.3 19.6k
Ours-unstop 52.3 41.3 44.1 36.8 43.6 +0.9 19.6k
Ours 52.3 41.6 44.4 37.3 43.9 +1.2 16.4k
Table 4. Experimental results (mAP@0.5) of SHIFT short-term CTTA task. We evaluate the performance by continually adapting the source model to the four conditions.
@@ 4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset. The results of the SHIFT short-term CTTA task are shown in Table 4. While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling and prompts learning achieves a 0.5 mAP improvement. Moreover, CTAOD which introduces a dynamic threshold strategy in AM shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
The experimental results of ACDC long-term CTTA task are summarized in Table 3. CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, CTAOD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines. While CTAOD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
@@ 4.4.3. Qualitative Results
As shown in Figure 3, we selectively compare the source model, MemCLR, and CTAOD in the final round adaptation to the brightness on the long-term Cityscapes-to-Cityscapes-C tasks. CTAOD guides the model in learning better feature representations while effectively mitigating forgetting in the long-term adaptation. Therefore, CTAOD assists the detector in distinguishing more foreground object categories and better locating them. More visualization analysis is presented in the appendix.
Mean-Teacher OCL AM DSR Mean Gain
1 $\checkmark$ 25.9 /
2 $\checkmark$ $\checkmark$ 26.7 +0.8
3 $\checkmark$ $\checkmark$ 27.3 +1.4
4 $\checkmark$ $\checkmark$ 27.4 +1.5
5 $\checkmark$ $\checkmark$ $\checkmark$ 28.6 +2.7
6 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29.2 +3.3
@@ 4.4.4. Ablation Study
We conduct an ablation study to empirically assess the impact of the OCL, AM, and DSR. As shown in Table 5, the integration of the mean-teacher architecture elevates the performance to 25.9 mAP. Building upon the mean-teacher foundation, the proposed OCL, AM, and DSR achieve an additional 1.5, 1.4, and 0.8 mAP improvement individually. Moreover, the combination of the OCL and AM modules raises the performance to 28.6 mAP. Furthermore, the inclusion of the DSR mechanism brings a 0.6 mAP boost in performance, culminating in a comprehensive model performance of 29.2 mAP. More fine-grained comparisons with AM and DSR are provided in the appendix.
@@ Appendix A Appendix Overview
In this appendix, we provide additional details and analysis. In the first section, we provide implementation details on CTAOD and other baseline methodologies. In the second section C, we furnish complete results for the quantitative research and additional visualization examples for the qualitative research. In the third section, we further discuss the continual learning. In the last section E, we discuss the limitations and prospective directions of our research.
@@ Appendix B Additional Implementation Details
Following (VS et al., 2023b), we trained a Faster R-CNN (Ren et al., 2015) with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) on the training set of Cityscapes (Cordts et al., 2016) as the source-trained model of Cityscapes. We utilized an SGD optimizer with a learning rate of 0.001, a momentum of 0.9, and a batch size of 1 and 70k iterations. Similarly, we trained another source model on the clear condition of the SHIFT (Sun et al., 2022) with 30 epochs. We will specify the implementation details of the CTAOD and other baseline methods in the following subsections.
@@ B.1. Implememtation of CTAOD
Detailed hyper-parameters of CTAOD for each benchmark are listed in Table 7. Furthermore, we present a pseudo-code outline of our CTAOD framework in Algorithm 1. The three key components of CTAOD are highlighted, i.e., Object-level Contrastive Learning, Adaptive Monitoring, and Data-driven Stochastic Restoration.
@@ B.2. Implememtation of Baseline Methods
The CoTTA (Wang et al., 2022) and SVDP (Yang et al., 2024) are originally designed for the classification or segmentation tasks. We implement them based on their public code, with some modifications to fit into the objection detection. Specifically, we implement the Cotta method by constructing a mean-teacher framework based on Faster R-CNN, which incorporates a stochastic restoration mechanism. Similarly, we implement SVDP with visual prompts on the area of the bounding box in pseudo-labels based on mean-teacher framework. Further, the MemCLR (VS et al., 2023b) and IRG (VS et al., 2023a) are source-free domain adaptation methods for object detection. Therefore, we implement the MemCLR and IRG based on their public code, with little modifications to follow the CTTA setting (i.e., updating the model parameters during evaluation). For a fair comparison, we set the same momentum $\alpha$ of the Exponential Moving Average (EMA) for CoTTA, SVDP, MemCLR, and IRG as CTAOD, all of which apply the mean-teacher framework (Tarvainen and Valpola, 2017). Other hyper-parameters are retained as specified in their public code.